OpenStudy (anonymous):

Determine if the graph is symmetric about the x-axis, the y-axis, or the origin. r = -3 - 2 cos θ

OpenStudy (anonymous):

@mathstudent55 @monsterfox1997 @loser66 @TheSmartOne @MD175

OpenStudy (anonymous):

@Rainbow_dashie

OpenStudy (anonymous):

@Michele_Laino @luckycoins888

OpenStudy (luckycoins888):

?

OpenStudy (michele_laino):

Please try to use cartesian coordinates

OpenStudy (luckycoins888):

yes

OpenStudy (michele_laino):

hint: use this transformation equations: \[\cos \theta =\frac{ x }{ r },\quad r ^{2}=x ^{2}+y ^{2}\]

OpenStudy (anonymous):

ok..... @Michele_Laino

OpenStudy (anonymous):

i tried and got stuck

OpenStudy (michele_laino):

Please, note that with the trasformation above I re-write your equation, as below: \[r=-3-2\frac{ x }{ r }\] so: \[r ^{2}=-3r-2x\] then: \[x ^{2}+y ^{2}=-3\sqrt{x ^{2}+y ^{2}}-2x\]